# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v2

@@ Abstract.
We formulate a precise connection between the new Drinfeld presentation of a quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$ and the new Drinfeld presentation of affine coideal subalgebras of split type recently discovered by Lu and Wang. In particular, we establish a “factorization formula”, expressing the commuting “affine Cartan”-type operators $\Theta_{i,k}$ in a coideal subalgebra in terms of the corresponding Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$. We study the spectra of these operators on finite dimensional representations, and describe them in terms of rational functions with an extra symmetry. These results can be seen as the starting point of a $q$-character theory for affine Kac–Moody quantum symmetric pairs. Additionally, we prove a compatibility result linking Lusztig’s and the Lu-Wang-Zhang braid group actions.
@@ 1.1. A trinity of presentations
The generalization of quantum symmetric pairs to the Kac–Moody case was carried out by Kolb [Kol14]. In particular, he provided a description of coideal subalgebras by generators and Serre-like relations. The RTT presentation of coideal subalgebras was constructed in special cases by Molev, Ragoucy and Sorba [MRS03], and was preceded by earlier work on twisted Yangians [Ols92]. On the other hand, the Drinfeld-type presentation of affine quantum symmetric pairs is a much more recent development. The first steps in this direction were taken in [BK20], where $q$-root vectors and some relations between them were computed in one of the rank $1$ cases (the $q$-Onsager algebra). A full Drinfeld-type presentation, for all simply laced split Satake diagrams, was later constructed by Lu and Wang [LW21], using a braid group action stemming from a Hall algebra realization of coideal subalgebras [LW23]. A generalization to non-simply laced Satake diagrams was obtained by Zhang [Zha22].
@@ 1.5.1. Factorization and coproduct theorems
Of particular importance in the Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ is the (almost) commutative subalgebra generated by the elements $\phi_{i,k}^{\pm}$. The spectra of these operators are used to define Drinfeld polynomials and $q$-characters. In a coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, there exists an analogous commutative subalgebra generated by elements $\Theta_{i,k}$. Our key result, which we call the factorization theorem, enables us to express the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ in terms of shifted series $\boldsymbol{\phi}_{i}(z)$, modulo a variant of the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details).
@@
(1.1) $\boldsymbol{\grave{\Theta}}_{i}(z)\equiv\boldsymbol{\phi}_{i}^{-}(-q^{-2}z^{-1})\boldsymbol{\phi}_{i}^{+}(-Cq^{-2}z){\mod U^{i}_{+}}[\negthinspace[z]\negthinspace].$
@@
$\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{(U^{i}_{-}\otimes U^{i}_{+}+U^{i}_{+}\otimes U^{i}_{-})}[\negthinspace[z]\negthinspace].$
@@ 1.5.3. Compatibility of braid group actions
It should be emphasized that the braid group action used to construct the Drinfeld presentation for coideal subalgebras in [LW21] is different from the usual Lusztig braid group action used to construct the Drinfeld presentation for quantum affine algebras. In particular, the former is not a restriction of the latter. Nevertheless, it turns out that, remarkably, the two actions are, roughly speaking, compatible modulo the “Drinfeld positive half” $U_{+}$ of $U_{q}\widehat{{\mathfrak{g}}}$. Not only is this a key ingredient allowing us to generalize Theorem 1.1 from the $q$-Onsager algebra to higher rank coideal subalgebras, but also a very interesting fact on its own. Together with Theorems 1.1 and 1.2, it suggests a general philosophy that although various structures on coideal subalgebras and quantum affine algebras are distinct, they are in fact compatible modulo $U_{+}$.
@@
Here $U_{\neq i,+}$ is the subalgebra spanned by elements of positive $Q_{i}$-degree, where $Q_{i}$ is the root sublattice of $Q$ spanned by all positive roots except $\alpha_{i}$. Theorem 1.4 is proven in §9 (see Theorem 9.7). For the details of the notation, see §8.
@@ 1.6.5. Factorization of the universal K-matrix
Originally, the notion of a $q$-character for quantum affine algebras was formulated in terms of the universal R-matrix [FR99]. Its connection to spectra of the operators $\phi_{i,k}^{\pm}$ was established using the Levendorsky–Soibelman factorization formula [KT94, LSS93, Dam98]. In the QSP case, we are proceeding in the opposite direction, i.e., we propose to define $q$-characters in terms of the spectra of the operators $\Theta_{i,k}$. We expect that an equivalent formulation in terms of the universal K-matrix from [AV22b] is also possible. In particular, there should exist an analogue of the Levendorsky–Soibelman formula for the universal K-matrix.
@@ Acknowledgements
I would like to express my gratitude for all the support I have received. I am particularly indebted to Andrea Appel, who has pointed out to me the importance of rationality properties and one dimensional representations, as well as to Jian-Rong Li for the careful reading of the manuscript and our ongoing collaboration on a related project. I would also like to thank Paul Terwilliger for insightful discussions about the $q$-Onsager algebra, Alexander Molev for correspondence about twisted Yangians, as well as Daniel Orr and Mark Shimozono for providing a reference on the combinatorics of the extended affine Weyl group.
@@ 2.2. The quantum loop algebra
The Serre and Drinfeld presentations are related by the following isomorphism:
$K\mapsto K,\quad e_{1}^{\pm}\mapsto x^{\pm}_{0},\quad e_{0}^{+}\mapsto x^{-}_{1}K^{-1},\quad e_{0}^{-}\mapsto Kx^{+}_{-1}.$
@@ 2.5. Coideal structures
(2.21) $A_{0}\mapsto x^{-}_{0}-c_{1}q^{2}K^{-1}x^{+}_{0}+s_{1}K^{-1},\quad A_{-1}\mapsto q^{-2}c_{0}^{-1}(Kx^{+}_{-1}-c_{0}x^{-}_{1}+s_{0}K).$
@@ 4.1. Auxiliary relations
$x^{+}_{k}\boldsymbol{\phi}(-q^{-2}z^{-1})=q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}+(q^{2}-q^{6})z\boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
$[x^{+}_{k},\boldsymbol{\phi}(-q^{-2}z^{-1})]=\boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k}.$
@@
(4.8) $x^{+}_{k}\boldsymbol{\phi}(-q^{-2}z^{-1})=$ $\ q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}+(q^{4}-1)\sum_{r\geq 0}\sum_{s=1}^{r}(-q^{2})^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s}z^{r}$
(4.9) $=$ $\ q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}$
(4.10) $+(q^{2}-q^{-2})(-q^{4})^{k}z^{k}\sum_{r\geq 0}\sum_{s=1}^{r}(-q^{2})^{r-s}\phi_{s-r}(-q^{4})^{s-k}x^{+}_{k-s}z^{r-k}$
(4.11) $=$ $\ q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}$
(4.12) $+(q^{2}-q^{-2})(-q^{4})^{k}z^{k}\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{x}^{+}_{\leq k-1}(-q^{-4}z^{-1})$
(4.13) $=$ $\ q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}+(q^{2}-q^{6})z\boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
(4.14) $[x^{+}_{k},\boldsymbol{\phi}(-q^{-2}z^{-1})]=$ $\ (q^{2}-1)\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{k}$
(4.15) $+(q^{2}-q^{6})z\boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}$
(4.16) $=$ $\ \boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}\big{(}(q^{2}-1)x^{+}_{k}+(q^{2}-q^{4})x^{+}_{k-1}z\big{)}$
(4.17) $=$ $\ \boldsymbol{\phi}(-q^{-2}z^{-1})(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k},$
@@
$(u-q^{-2}w)\boldsymbol{\phi}(u^{-1})\boldsymbol{x}^{-}(w^{-1})=(q^{-2}u-w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}(u^{-1}).$
Substituting $u^{-1}=-q^{-2}z^{-1}$ and multiplying both sides by $-1$, we get
$(q^{2}z+q^{-2}w)\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{x}^{-}(w^{-1})=(z+w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}(-q^{-2}z^{-1}).$
@@ 4.2. Basic implications of the Lu–Wang relations
(4.28) $=$ $\ [q^{2}(-c_{0}c_{1}\overline{h}_{1}+\overline{h}_{-1}),(-q^{2})^{r-1}x^{-}_{-r+1}]+C(-q^{2})^{r-2}x^{-}_{-r+2}$
(4.29) $=$ $\ (-q^{2})^{r}(x^{-}_{-r}-c_{0}c_{1}x^{-}_{-r+2})+c_{0}c_{1}(-q^{2})^{r}x^{-}_{-r+2}$
(4.30) $=$ $\ (-q^{2})^{r}x^{-}_{-r}$
for $r\geq 1$. A similar argument using the relation $A_{r}=(q^{4}c_{0}c_{1})^{-1}(A_{r+2}-[\overline{H}_{1},A_{r+1}])$ and downward induction establishes (4.26) for $r\leq-2$. ∎
@@ 4.3. $\boldsymbol{A}^{(1)}_{+}(z)$ calculation
(4.32) $\Omega=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)(1+c_{0}c_{1}q^{2}\operatorname{ad}_{\overline{h}_{1}}z).$
@@
(4.33) $\boldsymbol{A}^{(1)}_{+}(z)=$ $\ c_{1}\boldsymbol{\phi}(-q^{-2}z^{-1})\Omega^{-1}\big{(}-(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(-q^{-4}z^{-1})-q^{2}x^{+}_{0}+q^{6}x^{+}_{-1}z\big{)}.$
@@
(4.37) $\Omega\boldsymbol{A}^{(1)}_{+}(z)-z\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=q^{2}c_{1}(Kx^{+}_{-1}z-K^{-1}x^{+}_{0}).$
@@
(4.38) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ \frac{c_{1}}{[2]}\sum_{r\geq 0}(-q^{2})^{r}\big{[}[x^{+}_{0},x^{+}_{-1}]_{q^{6}},x^{-}_{-r}\big{]}z^{r}.$
@@
(4.46) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ -c_{1}q^{2}Kx^{+}_{-1}+c_{1}\sum_{r\geq 0}(-q^{2})^{r}(q^{4}\phi_{-r-1}x^{+}_{0}+q^{6}\phi_{-r}x^{+}_{-1})z^{r}$
(4.47) $\ +c_{1}(q^{4}-1)\sum_{r\geq 0}\sum_{s=2}^{r+1}q^{2s}(-q^{2})^{r}\phi_{s-r-1}x^{+}_{-s}z^{r}.$
(4.48) $=$ $\ -c_{1}q^{2}Kx^{+}_{-1}+c_{1}q^{2}K^{-1}x^{+}_{0}z^{-1}-c_{1}q^{2}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{0}z^{-1}$
(4.49) $\ +c_{1}q^{6}\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{-1}-c_{1}(q^{4}-1)q^{-2}\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{x}^{+}_{\leq-2}(-q^{-4}z^{-1})z^{-1}.$
@@
(4.50) $\Omega\boldsymbol{A}^{(1)}_{+}(z)=$ $\ c_{1}\boldsymbol{\phi}(-q^{-2}z^{-1})\big{(}-(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(-q^{-4}z^{-1})-q^{2}x^{+}_{0}+q^{6}x^{+}_{-1}z\big{)}.$
Dividing both sides by $\Omega$ and using the fact that $\Omega$ commutes with $\boldsymbol{\phi}(-q^{-2}z^{-1})$ completes the proof. ∎
@@
(4.51) $\mathbb{A}=c_{1}\Omega^{-1}\big{(}-(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(-q^{-4}z^{-1})-q^{2}x^{+}_{0}+q^{6}x^{+}_{-1}z\big{)},$
so that $\boldsymbol{A}^{(1)}_{+}(z)=\boldsymbol{\phi}(-q^{-2}z^{-1})\mathbb{A}$.

### v3

@@ Abstract.
We formulate a precise connection between the new Drinfeld presentation of a quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$ and the new Drinfeld presentation of affine coideal subalgebras of split type recently discovered by Lu and Wang. In particular, we establish a “factorization formula”, expressing the commuting “Drinfeld–Cartan”-type operators $\Theta_{i,k}$ in a coideal subalgebra in terms of the corresponding Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$. We study the spectra of these operators on finite dimensional representations, and describe them in terms of rational functions with an extra symmetry. These results can be seen as the starting point of a $q$-character theory for affine Kac–Moody quantum symmetric pairs. Additionally, we prove a compatibility result linking Lusztig’s and the Lu-Wang-Zhang braid group actions.
@@ 1.1. A trinity of presentations
The generalization of quantum symmetric pairs to the Kac–Moody case was carried out by Kolb [Kol14]. In particular, he provided a description of coideal subalgebras by generators and Serre-like relations. The RTT presentation of coideal subalgebras was constructed in special cases by Molev, Ragoucy and Sorba [MRS03], and was preceded by earlier work on twisted Yangians [Ols92]. On the other hand, the Drinfeld-type presentation of affine quantum symmetric pairs is a much more recent development. The first steps in this direction were taken by Baseilhac and Kolb in [BK20], where $q$-root vectors and the relations between them were computed in one of the rank $1$ cases (the $q$-Onsager algebra). Their approach was inspired by Damiani’s work for $\widehat{\mathfrak{sl}}_{2}$ [Dam93], and involved the construction of a new braid group action on the $q$-Onsager algebra, using a computer-based proof. A computer-free proof of the existence of this braid group action was given independently by Terwilliger in [Ter18]. A Drinfeld-type presentation for all simply laced split Satake diagrams was later constructed by Lu and Wang [LW21], using a more general braid group action stemming from a Hall algebra realization of coideal subalgebras [LW23]. A generalization to non-simply laced Satake diagrams was obtained by Zhang [Zha22].
@@ 1.5.1. Factorization and coproduct theorems
Of particular importance in the Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ is the (almost) commutative subalgebra generated by the elements $\phi_{i,k}^{\pm}$. The spectra of these operators are used to define Drinfeld polynomials and $q$-characters. In a coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, there exists an analogous commutative subalgebra generated by elements $\Theta_{i,k}$. Our key result, which we call the factorization theorem, enables us to express the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ in terms of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details).
@@
(1.1) $\boldsymbol{\grave{\Theta}}_{i}(z)\equiv\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
@@
$\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{(U_{-}\otimes U_{+}+U_{+}\otimes U_{-})}[\negthinspace[z]\negthinspace].$
@@ 1.5.3. Compatibility of braid group actions
It should be emphasized that the braid group action used to construct the Drinfeld presentation for coideal subalgebras in [BK20, LW21] is different from the usual Lusztig braid group action used to construct the Drinfeld presentation for quantum affine algebras. In particular, the former is not a restriction of the latter. Nevertheless, it turns out that, remarkably, the two actions are, roughly speaking, compatible modulo the “Drinfeld positive half” $U_{+}$ of $U_{q}\widehat{{\mathfrak{g}}}$. Not only is this a key ingredient allowing us to generalize Theorem 1.1 from the $q$-Onsager algebra to higher rank coideal subalgebras, but also a very interesting fact on its own. Together with Theorems 1.1 and 1.2, it suggests a general philosophy that although various structures on coideal subalgebras and quantum affine algebras are distinct, they are in fact compatible modulo $U_{+}$.
@@
Here $U_{\neq i,+}$ is the subalgebra spanned by elements of positive $Q_{i}$-degree, where $Q_{i}$ is the root sublattice of $Q$ spanned by all positive roots except $\alpha_{i}$. Theorem 1.4 is proven in §9 (see Corollary 9.7). For the details of the notation, see §8.
@@ 1.6.5. Factorization of the universal K-matrix
Originally, the notion of a $q$-character for quantum affine algebras was formulated in terms of the universal R-matrix [FR99]. Its connection to spectra of the operators $\phi_{i,k}^{\pm}$ was established using the Levendorsky–Soibelman factorization formula [KT94, LSS93, Dam98]. In the QSP case, we are proceeding in the opposite direction, i.e., we propose to define $q$-characters in terms of the spectra of the operators $\Theta_{i,k}$. We expect that an equivalent formulation in terms of the universal K-matrix from [AV22b] is also possible. In particular, there should exist an analogue of the Levendorsky–Soibelman formula for the universal K-matrix. In the finite case, similar factorization results were already established in [WZ23].
@@ Acknowledgements
I would like to express my gratitude for all the support I have received. I am particularly indebted to Andrea Appel, who has pointed out to me the importance of rationality properties and one dimensional representations, as well as to Jian-Rong Li for the careful reading of the manuscript and our ongoing collaboration on a related project. I would also like to thank Paul Terwilliger for insightful discussions about the $q$-Onsager algebra, Alexander Molev for correspondence about twisted Yangians, Daniel Orr and Mark Shimozono for providing a reference on the combinatorics of the extended affine Weyl group, as well as Pascal Baseilhac, Weiqiang Wang and Weinan Zhang for comments on the first version of this paper.
@@ 2.2. The quantum loop algebra
The Serre and Drinfeld presentations are related by the following isomorphism ([Bec94, Proposition 3.7]):
$K\mapsto K,\quad e_{1}^{\pm}\mapsto x^{\pm}_{0},\quad e_{0}^{+}\mapsto-K^{-1}x^{-}_{1},\quad e_{0}^{-}\mapsto-x^{+}_{-1}K.$
@@ 2.5. Coideal structures
(2.21) $A_{0}\mapsto x^{-}_{0}-c_{1}q^{2}K^{-1}x^{+}_{0}+s_{1}K^{-1},\quad A_{-1}\mapsto-q^{-4}c_{0}^{-1}Kx^{+}_{-1}+x^{-}_{1}+q^{-2}c_{0}^{-1}s_{0}K.$
@@ 4.1. Auxiliary relations
$x^{+}_{k}\boldsymbol{\phi}(z^{-1})=q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
$[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k}.$
@@
(4.8) $x^{+}_{k}\boldsymbol{\phi}(z^{-1})=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)\sum_{r\geq 0}\sum_{s=1}^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s}z^{r}$
(4.9) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\sum_{r\geq 0}\sum_{s=1}^{r}\phi_{s-r}q^{2(s-k)}x^{+}_{k-s}z^{r-k}$
(4.10) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq k-1}(q^{-2}z^{-1})$
(4.11) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
(4.12) $[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=$ $\ (q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}$
(4.13) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(x^{+}_{k}+x^{+}_{k-1}z)$
(4.14) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k},$
@@
$(z-q^{-2}w)\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{-}(w^{-1})=(q^{-2}z-w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}(z^{-1}).$
@@ 4.2. Basic implications of the Lu–Wang relations
(4.25) $=$ $\ [C\overline{h}_{1}-\overline{h}_{-1}),x^{-}_{-r+1}]+Cx^{-}_{-r+2}$
(4.26) $=$ $\ -Cx^{-}_{-r+2})+x^{-}_{-r}+Cx^{-}_{-r+2}$
(4.27) $=$ $\ x^{-}_{-r}$
for $r\geq 1$. A similar argument using the relation $A_{r}=C^{-1}(A_{r+2}-[\overline{H}_{1},A_{r+1}])$ and downward induction establishes (4.23) for $r\leq-2$. ∎
@@ 4.3. $\boldsymbol{A}^{(1)}_{+}(z)$ calculation
(4.29) $\Omega=(1+\operatorname{ad}_{\overline{h}_{-1}}z)(1-C\operatorname{ad}_{\overline{h}_{1}}z).$
@@
(4.30) $\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
@@
(4.34) $\Omega\boldsymbol{A}^{(1)}_{+}(z)-z\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=-c_{1}(Kx^{+}_{-1}z+q^{2}K^{-1}x^{+}_{0}).$
@@
(4.35) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ -q^{-2}\frac{c_{1}}{[2]}\sum_{r\geq 0}\big{[}[x^{+}_{0},x^{+}_{-1}]_{q^{6}},x^{-}_{-r}\big{]}z^{r}.$
@@
(4.43) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ c_{1}Kx^{+}_{-1}-q^{-2}c_{1}\sum_{r\geq 0}(q^{4}\phi_{-r-1}x^{+}_{0}+q^{6}\phi_{-r}x^{+}_{-1})z^{r}$
(4.44) $\ -c_{1}(q^{2}-q^{-2})\sum_{r\geq 0}\sum_{s=2}^{r+1}q^{2s}\phi_{s-r-1}x^{+}_{-s}z^{r}.$
(4.45) $=$ $\ c_{1}Kx^{+}_{-1}+c_{1}q^{2}K^{-1}x^{+}_{0}z^{-1}-c_{1}q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{0}z^{-1}$
(4.46) $\ -c_{1}q^{4}\boldsymbol{\phi}(z^{-1})x^{+}_{-1}-c_{1}(q^{2}-q^{-2})\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})z^{-1}.$
@@
(4.47) $\Omega\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
Dividing both sides by $\Omega$ and using the fact that $\Omega$ commutes with $\boldsymbol{\phi}(z^{-1})$ completes the proof. ∎
@@
(4.48) $\mathbb{A}=-c_{1}\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)},$
so that $\boldsymbol{A}^{(1)}_{+}(z)=\boldsymbol{\phi}(z^{-1})\mathbb{A}$.